# Probing the Turbulent Corona and Heliosphere Using Radio Spectral Imaging Observation during the Solar Conjunction of Crab Nebula

### v1

@@ 1 Introduction
With spatially resolved radio observation from radio telescope arrays, it is possible to observe the angular broadening and orientation of the broadening to derive the scattering properties of the plasma density fluctuation and its anisotropic degree. Early studies with the Very Large Array (Armstrong et al., 1990; Grall et al., 1997) have demonstrated that the anisotropic ratio of the source can be derived by forward fitting the visibility with free parameters, including the axial ratio and the orientation angle at the frequencies of 1.5 and 4.8 GHz. The Crab Nebula (Tau-A), a bright and relatively compact source in low frequencies (200MHz), is particularly suitable for the study of angular broadening. For example, Ramesh et al. (2001) reported an increase in the amplitude of the angular broadening of Tau-A when a CME crossed the line of sight, indicating there is enhanced density fluctuation within the CME. Raja et al. (2016, 2017, 2019) discussed the density turbulence and proton heating rate based on the observation of Tau-A’s solar conjunction in 2012-2013.
@@
Despite the long-term interest in this area, there remains a lack of low-frequency wideband observations that explore the frequency-dependent behavior of anisotropic angular broadening. As shown by Kontar et al. (2023), multiple factors, including the electron density, the amplitude of density fluctuations, and the degree of anisotropy in the fluctuations, govern angular broadening. These parameters are strongly coupled, making the interpretation of angular broadening complex. Observations at a single frequency can provide limited quantitative diagnostics, while multi-frequency imaging is essential for more comprehensive and robust constraints.. By comparing the parametric simulation and the multi-frequency observation, one may place better constraints on the multiple factors above and as well as providing a diagnosis on the plasma property. Additionally, high-sensitivity and high-dynamic-range observations are needed to investigate features associated with solar approaches at low frequencies.
@@
The brightness of Tau-A at low frequencies () is stable, making it a widely used calibration source. Its flux spectrum has been extensively measured and is well known (De Gasperin et al., 2020). However, when Tau-A’s line of sight is the Sun, the plasma with a higher and presumably more fluctuating electron density in the solar vicinity causes increased scattering and absorption of the radio waves (Kontar et al., 2019, 2023). Consequently, the observed total flux density is reduced. As the scattering and absorption effect is stronger for lower frequencies, the effect of reduced flux is more pronounced. The observation of the flux density of the source and its frequency dependency can thus be used to estimate the absorption parameters, as well as the distribution and fluctiations of density in the outer corona and inner heliosphere.
@@
In addition to probing the turbulence properties of the general solar corona, it is also of interest to detect such properties within specific coronal structures, such as streamers. Solar streamers are fundamental components of the coronal architecture (Brueckner et al., 1995). However, measuring the extended density distribution of streamers, particularly in the outer corona, remains challenging. Previous efforts, such as Decraemer et al. (2019), have combined white-light coronagraph observations with forward modeling to infer the density distribution of streamers. In this context, the density variation of the streamer can be embedded into the observed source when its line of sight passes through the streamer, This potential capability offers a novel viewing angle to investigate streamer structures. Despite its potential, such use of radio source substructures to study large-scale coronal features has not been systematically explored before.
@@ 2 Observation
Apply bandpass solutions from nighttime calibrator observation,
@@
Imaging step to convert visibility into the flux density distribution image by gridding and CLEAN process to de-convolution the point-spread-function (PSF) from the resulting image.
@@
Finally, cut out the region of interest and do source fitting to get the source brightness, size, and shape. Prepare for further analysis.
@@ 2.1 Angular broadening and orientation
The deconvolved source shape provides a more accurate estimate of the intrinsic source properties by minimizing the influence of the synthesized beam on the observed image. Figure 2 shows the fitted result of the deconvolved 2D Gaussian model. Column (a) shows the orientation, defined as the angular difference between the source’s minor axis () and the radial direction connecting the Crab to the center of the Sun (). Column (b) displays the full width at half maximum (FWHM) of the source. Column (c) shows the axial ratio, defined by the ratio of major to minor axis. Summary statistics and fitting results are also shown in Table 1.
@@
When the source displays a significant eccentricity, the orientation (as shown in panel (a) in Figure 2) of the source’s minor axis aligns very well with the radial direction of the Sun (the average deviation angle is within 8 degrees). This alignment indicates that turbulence-induced scattering is stronger in the tangential direction—perpendicular to the solar radial direction.
@@ 2.2 Source deformation near solar transit
In the model, , is the polar coordinate element, is the rms width in the radial direction, and is the rms angular width of the source measured from the source center (, ). The radial and tangential direction are defined with respect to the coordinate center of the Sun obtained from the image with the method introduced in.
@@
On June 13th and 15th, when Tau A was during its ingress and egress of the solar transit at small projected heliocentric distances (6.4 and 5.6 , respectively), the source exhibits distinct substructures at lower frequency bands (36 MHz and 55 MHz). Despite this, its overall arc-shaped brightness distribution can still be characterized using the anisotropic polar Gaussian model. The brightness distribution along the angular direction (Figure 5 panels (a 1-4) and (b 1-4)) indicates that there are more substructures in the source of June 13th (ingress) than that of June 15th (egress): on June 13th at 36 MHz, the source exhibits three prominent peaks, and at 55 MHz the angular distribution shows two well-defined peaks. A notable feature is the anti-correlation between the angular distributions of the white-light and radio emissions (panels (a3) and (a0)). Specifically, regions with enhanced white-light brightness—indicative of higher plasma density—correspond to regions of diminished radio flux. This suggests that the radio wave is absorbed or re-distributed by the higher density streamers. For the source observed on June 15th, however, the angular distribution presents much less prominent substructures. Also, the source size in the tangential direction on the 13th is larger than on the 15th. According to the white light coronagraph data, a clear difference during the ingress and egress at similar heliocentric distances lies in that the source’s line of sight on June 13th passes through a more pronounced streamer region compared to that on June 15th, as shown in Figures 5(a) and (b). The observed sub-structures in the source morphologies suggest that, on top of the scattering effects induced by the overall coronal plasma, they may carry additional information about density fluctuations associated with smaller-scale coronal structures, such as streamers, at spatial scales of 0.1–1∘.
@@ 3 Scattering and anisotropic ratio
The angular broadening of a distant point source with its line-of-sight passing through the turbulent corona or inner heliosphere has been modeled by ray tracing simulations (Kontar et al., 2019, 2023). The model can be used to emulate the scattering process along the line of sight and estimate the observed angular size of the broadened source. There is a key factor in the simulation: the anisotropic ratio (), which represents the ratio of the wavenumber of the density fluctuation in the parallel or perpendicular direction. It was reported as in the range of 0.25-0.5 using various types of observations. In this work, to compare with the observed source shape, we calculate the source size of a remote point source according to Equations (10-11) in Kontar et al. (2023), using an assumed value of . The model-predicted angular FWHM size, shown in Figure 7 as black dashed and dotted lines for the major and minor axes, respectively, has a constant source axial ratio from 2 to 40 solar radii.
From observation, we can measure the size of the major and minor axis during the ingress and egress period. The fitted result is:
@@
However, the observed axial ratio exhibits a systematic deviation from the model predictions: it is generally larger than the model at smaller heliocentric distances and smaller than the model at larger distances (as shown in Figure 7(e)). This trend suggests that the plasma density fluctuations closer to the Sun are more anisotropic than assumed in the model with a constant , while those further away are less anisotropic than . At approximately , the observed source shape is consistent with the model prediction using . Thus, our results indicate that there is a radial variation in anisotropy, with stronger anisotropy in the inner corona and weaker anisotropy in outer corona.
@@ 4 Discussion and Conclusion
The observed anisotropic feature of the broadened source, in general, agrees with the predictions from the stochastic wave scattering model (Kontar et al., 2023). The measured value of axial ratio () ranges from 1.7 to 2.5, which is consistent with the scattering model (Kontar et al., 2023) predicted a fix value of 2.0 (derived with ) in the range of 6-40 solar radius. The shape axial ratio of the source has a trend of increasing with the decrease of the separation angle, indicating more anisotropic density fluctuations for the coronal medium near the Sun.
@@
For the first time, sub-structures within the angular broadened source were observed. These substructures appeared when the source’s line of sight intersected strong streamer regions in the corona. We found anti-correlation between the brightness angular distribution of white-light and the substructure.
@@
The observed features in this study include the frequency and distance dependence of angular broadening, arc-shaped source morphology, substructures within the broadened sources, and brightness reduction near the Sun. These features are strongly linked to coronal structures across multiple spatial scales. The angular broadening itself arises from small-scale density fluctuations along the line of sight, with its shape reflecting the amplitude and anisotropic ratio of the turbulence. Meanwhile, the radial trend in broadening and brightness attenuation is related to large-scale variations in the background coronal density and fluctuation amplitude. Meso-scale structures such as coronal streamers further modulate the scattering: when the line of sight intersects these dense, structured regions, the resulting broadened source exhibits arc-like deformation and internal substructures. These signatures carry the imprint of streamer-scale density gradients and inhomogeneities. Thus, angular broadening is a good indicator of the density variation of the coroal across different scales. Forward modeling of radio-wave scattering through anisotropic turbulence, as demonstrated in Kontar et al. (2023), provides a theoretical framework to interpret these observables. With future observations of additional compact sources near the Sun and expanded multi-frequency coverage, combined with advanced wave propagation simulations, we can place tighter constraints on key coronal parameters—including the spatial distribution of density fluctuations, streamer profiles, and the variation of anisotropy with heliocentric distance.

### v2

@@ 1 Introduction
With spatially resolved radio observation from radio telescope arrays, it is possible to observe the angular broadening and orientation of the broadening to derive the scattering properties of the plasma density fluctuation and its anisotropic degree. Early studies with the Very Large Array (Armstrong et al., 1990; Grall et al., 1997) have demonstrated that the anisotropic ratio of the source can be derived by forward fitting the visibility with free parameters, including the axial ratio and the orientation angle at the frequencies of 1.5 and 4.8 GHz. The Crab Nebula (Tau-A), a bright and relatively compact source in low frequencies (200MHz), is particularly suitable for the study of angular broadening. The angular broadening phenomena were first reported by Hewish (1955, 1958), who measured the size of Tau A with an interferometer, at wavelengths of 3.7 m and 7.9 m. Ramesh et al. (2001) reported an increase in the amplitude of the angular broadening of Tau-A when a CME crossed the line of sight, indicating there is enhanced density fluctuation within the CME. Sasikumar Raja et al. (2016, 2017, 2019) discussed the density turbulence and proton heating rate based on the observation of Tau-A’s solar conjunction in 2012-2013.
@@
Despite the long-term interest in this area, there remains a lack of low-frequency wideband observations that explore the frequency-dependent behavior of anisotropic angular broadening. As shown by Kontar et al. (2023), multiple factors, including the electron density, the amplitude of density fluctuations, and the degree of anisotropy in the fluctuations, govern angular broadening. These parameters are strongly coupled, making the interpretation of angular broadening complex. Observations at a single frequency can provide limited quantitative diagnostics, while multi-frequency imaging is essential for more comprehensive and robust constraints. By comparing the parametric simulation and the multi-frequency observation, one may place better constraints on the multiple factors above and as well as providing a diagnosis on the plasma property. Additionally, high-sensitivity and high-dynamic-range observations are needed to investigate features associated with solar approaches at low frequencies.
@@
Tau-A at low frequencies () is bright and stable, making it a widely used calibration source. Its flux spectrum has been extensively measured and is well known (De Gasperin et al., 2020). However, when Tau-A’s line of sight passes through the solar corona, the plasma with a higher and presumably more fluctuating electron density in the solar vicinity causes increased scattering and absorption of the radio waves (Kontar et al., 2019, 2023). Consequently, the observed total flux density is reduced. As the scattering and absorption effect is stronger for lower frequencies, the effect of reduced flux is more pronounced. The observation of the flux density of the source and its frequency dependency can thus be used to estimate the absorption parameters, as well as the distribution and fluctuations of density in the outer corona and inner heliosphere.
@@
In addition to probing the turbulence properties of the general solar corona, it is also of interest to detect such properties within specific coronal structures, such as streamers. Solar streamers are fundamental components of the coronal architecture (Brueckner et al., 1995). However, measuring the extended density distribution of streamers, particularly in the outer corona, remains challenging. Previous efforts, such as Decraemer et al. (2019), have combined white-light coronagraph observations with forward modeling to infer the density distribution of streamers. In this context, the density variation of the streamer can be embedded into the observed source when its line of sight passes through the streamer. This potential capability offers a novel viewing angle to investigate streamer structures. Despite its potential, such use of radio source substructures to study large-scale coronal features has not been systematically explored before.
@@ 2 Observation
Apply bandpass solutions from the nighttime calibrator observation.
@@
Imaging step to convert visibility into the flux density distribution image by gridding and CLEAN process to deconvolve the point-spread-function (PSF) from the resulting image.
@@
Finally, cutting out the region of interest and performing source fitting to get the source brightness, size, and shape. Prepare for further analysis.
@@ 2.1 Angular broadening and orientation
The deconvolved source shape provides a more accurate estimate of the intrinsic source properties by minimizing the influence of the synthesized beam on the observed image. Figure 2 shows the fitted result of the deconvolved 2D Gaussian model. Column (a) shows the orientation, defined as the angular difference between the source’s minor axis () and the radial direction connecting the Crab to the center of the Sun (). Column (b) displays the full width at half maximum (FWHM) of the source. Column (c) shows the axial ratio, defined by the ratio of major to minor axis. Summary statistics and fitting results are also shown in Table 1. The intrinsic size of Tau A, as reported by De Gasperin et al. (2020), is 7.9 arcminutes with a bright compact central core at 50MHz. This size is indicated by a green circle in the upper-left corner of Figure 1. At higher frequencies ( MHz), the source size is comparable to the synthesized beam, and at lower frequencies ( MHz), it is smaller than the beam. Therefore, Tau A is treated as an unresolved source throughout this work.
@@
When the source displays a significant eccentricity, the orientation (as shown in panel (a) in Figure 2) of the source’s minor axis aligns very well with the radial direction of the Sun (the average deviation angle is within 8 degrees). This alignment indicates that turbulence-induced scattering is stronger in the tangential direction, perpendicular to the solar radial direction.
@@ 2.2 Source deformation near solar transit
In the model, , is the polar coordinate element, is the rms width in the radial direction, and is the rms angular width of the source measured from the source center (, ). The radial and tangential directions are defined with respect to the coordinate center of the Sun obtained from the image with the method introduced in.
@@
On June 13th and 15th, when Tau A was during its ingress and egress of the solar transit at small projected heliocentric distances (6.4 and 5.6 , respectively), the source exhibits distinct substructures at lower frequency bands (36 MHz and 55 MHz). Despite this, its overall arc-shaped brightness distribution can still be characterized using the anisotropic polar Gaussian model. The brightness distribution along the angular direction (Figure 5 panels (a 1-4) and (b 1-4)) indicates that there are more substructures in the source of June 13th (ingress) than that of June 15th (egress): on June 13th at 36 MHz, the source exhibits three prominent peaks, and at 55 MHz the angular distribution shows two well-defined peaks. A notable feature is the anti-correlation between the angular distributions of the white-light and radio emissions (panels (a3) and (a0)). Specifically, regions with enhanced white-light brightness—indicative of higher plasma density—correspond to regions of diminished radio flux. This suggests that the radio wave is absorbed or re-distributed by the higher-density streamers. For the source observed on June 15th, however, the angular distribution presents much less prominent substructures. Also, the source size in the tangential direction on the 13th is larger than on the 15th. According to the white light coronagraph data, a clear difference during the ingress and egress at similar heliocentric distances lies in that the source’s line of sight on June 13th passes through a more pronounced streamer region compared to that on June 15th, as shown in Figures 5(a) and (b). The observed sub-structures in the source morphologies suggest that, on top of the scattering effects induced by the overall coronal plasma, they may carry additional information about density fluctuations associated with smaller-scale coronal structures, such as streamers, at spatial scales of 0.1–1∘.
@@ 3 Scattering and anisotropic ratio
The angular broadening of a distant point source with its line-of-sight passing through the turbulent corona or inner heliosphere has been modeled by ray tracing simulations (Kontar et al., 2019, 2023). The model can be used to emulate the scattering process along the line of sight and estimate the observed angular size of the broadened source. There is a key factor in the simulation: the anisotropic ratio (), which represents the ratio of the wavenumber of the density fluctuation in the parallel or perpendicular direction. It was reported as in the range of 0.25-0.5 using various types of observations. In this work, to compare with the observed source shape, we derived the expected observed source shape with numerical simulation with the model in Kontar et al. (2023), using an assumed value of . The model-predicted angular FWHM size, shown in Figure 7 as black dashed and dotted lines for the major and minor axes, respectively, has a constant source axial ratio from 2 to 40 solar radii.
From observation, we can measure the size of the major and minor axes during the ingress and egress period. The fitted results are as follows, we note that the source shape from June 13 to June 15 is not included in the fitting for the consideration of different fitting schemes and occultation:
@@
However, the observed axial ratio exhibits a systematic deviation from the model predictions: it is generally larger than the model at smaller heliocentric distances and smaller than the model at larger distances (as shown in Figure 7(e)). This trend suggests that the plasma density fluctuations closer to the Sun are more anisotropic than assumed in the model with a constant , while those further away are less anisotropic than . At approximately , the observed source shape is consistent with the model prediction using . Thus, our results indicate that there is a radial variation in anisotropy, with stronger anisotropy in the inner corona and weaker anisotropy in the outer corona.
@@ 4 Discussion and Conclusion
The observed anisotropic feature of the broadened source, in general, agrees with the predictions from the stochastic wave scattering model (Kontar et al., 2023). The measured value of axial ratio () ranges from 1.7 to 2.5, which is consistent with the scattering model (Kontar et al., 2023) predicted a fixed value of 2.0 (derived with ) in the range of 6-40 solar radius. The shape axial ratio of the source has a trend of increasing with the decrease of the separation angle, indicating more anisotropic density fluctuations for the coronal medium near the Sun.
@@
For the first time, sub-structures within the angular broadened source were observed. These substructures appeared when the source’s line of sight intersected strong streamer regions in the corona. We found an anti-correlation between the brightness angular distribution of white-light and the substructure.
@@
The size and shape reported in this work are, in general, in good agreement with previous observations. The size from the early work by Hewish & Wyndham (1963) was measured with the fringe pattern of the interferometer at 38 MHz, with the assumption that the brightness profile follows a Gaussian distribution as a function of baseline length. The observation showed that the size distribution as a function of the heliocentric distance could be described by a power-law function with a form of , with a power-law index ranging from 1.30 to 2.24 in the range of 10–20 based on observations from 1959 to 1962. We have similar results in the range of 6–19 , with a power-law index ranging from to for for major and minor axes of ingress and egress, respectively. The axial ratio is difficult to determine directly from interferometric fringe patterns. In previous works (Hewish, 1958; Hewish et al., 1964), an assumed value of 2:1 was used. Sasikumar Raja et al. (2017) reported an axial ratio of approximately 2 with interferometric imaging observations, with larger values observed in the presence of streamers, though no systematic variation with heliocentric distance was noted. In this work, we also measure an axial ratio of , consistent with previous studies, but we additionally find a clear trend of increasing axial ratio as the line of sight approaches the Sun.
The observed features in this study include the frequency and distance dependence of angular broadening, arc-shaped source morphology, substructures within the broadened sources, and brightness reduction near the Sun. These features are strongly linked to coronal structures across multiple spatial scales. The angular broadening itself arises from small-scale density fluctuations along the line of sight, with its shape reflecting the amplitude and anisotropic ratio of the turbulence. Meanwhile, the radial trend in broadening and brightness attenuation is related to large-scale variations in the background coronal density and fluctuation amplitude. Meso-scale structures such as coronal streamers further modulate the scattering: when the line of sight intersects these dense, structured regions, the resulting broadened source exhibits arc-like deformation and internal substructures. These signatures carry the imprint of streamer-scale density gradients and inhomogeneities. Thus, angular broadening is a good indicator of the density variation of the coronal across different scales. Forward modeling of radio-wave scattering through anisotropic turbulence, as demonstrated in Kontar et al. (2023), provides a theoretical framework to interpret these observables. With future observations of additional compact sources near the Sun and expanded multi-frequency coverage, combined with advanced wave propagation simulations, we can place tighter constraints on key coronal parameters—including the spatial distribution of density fluctuations, streamer profiles, and the variation of anisotropy with heliocentric distance.